4) Given the following information what is the p-value of the test? Using the p-value method would you reject or not reject the null hypothesis at a significance level of 0.05? (8 pts)

           

a) The test statistic for a right-tailed test is z = 0.55

b) The test statistic for a two-tailed test is z = 1.95

c) The test statistic in a left-tailed test is z = -1.72.

d) The test statistic for a two-tailed test is z = -1.63.
Answer #1

If p - value will be less than 0.05 then we will reject the null hypothesis. Hence,

a) p - value = P(z > 0.55) = 0.2912 > 0.05

Do not reject Ho

b) p - value = 2*P(z > 1.95) = 2*0.0256 = 0.0512 > 0.05

Do not reject Ho

c) p - value = P(z < -1.72) = 0.0427 < 0.05

Reject Ho

d) p - value = 2*P(z < -1.63) = 2*0.0516 = 0.1032 > 0.05

Do not reject Ho